1. Assume an economy where spending for each sector is:

Household: C = 800 + 0.95Q

Business: I = 3000

Public: G = 4000, Tr = 7000, Tx = 1000 + 0.3Q

Foreign: X = 1700, Im = 200 + 0.165Q

Solve for a) Autonomous Spending, b) Spending Multiplier, c) Disposable Income, d) Consumption Expenditure, e) Household Savings,

f) Imports, g) Net Exports, h) Government Expenditure, i) Budget Deficit

2.) Continuing with the previous problem, what happens to the values in parts a-i if we increase Government Purchases by 1500?

3.) Continuing with 2, what happens if Investment Spending decreased by 500?

4.) Continuing with 3, what happens if exports also decrease by 500?
1
Expert's answer
2020-04-14T10:09:10-0400


1.Q=800+0.95Q+3000+4000+1700-(200+0.165Q)

Q=800+0.95q+3000+4000+1700-200-0.165Q

Q-0.785Q=9300

0.215Q=9300

Q=43 255.81

a)autonomous expenses=800

total autonomous expenses = autonomous consumption + autonomous investments +

government spending + autonomous net export =800+3000+4000+1500=9500.

b)MPS = 1-MPC = 1-0.95 = 0,05

"M=\frac{1}{MPS}=\frac{1}{0.05}=20"

c)Disposable Income=C+S-Tx=43 255.81-(1000+0.3Q)=43 255.81-1000-12976.74=29 279.07

C=800+0.95Q=800+0.95*43 255.81=41 893.02

S=-800+0.05Q=-800+0.05*43 255.81=1 362.79

d)C=800+0.95Q=800+0.95*43 255.81=41 893.02

e)S=-800+0.05Q=-800+0.05*43 255.81=1 362.79

f)Im = 200 + 0,165Q=200+0.165*43 255.81=7337.21

g)NX=X-Im=1700-200-0.165Q=1500-0.165Q=1500-7137.21=-5637.21

h)Government Expenditure=4000+7000=11 000

i)Budget Deficit =11 000-1000-0.3Q=-2976.74

2)

Q=800+0.95Q+3000+5500+1700-(200+0.165Q)

Q=800+0.95q+3000+5500+1700-200-0.165Q

Q-0.785Q=10 800

Q=50 232.56

a)autonomous expenses=800

total autonomous expenses = autonomous consumption + autonomous investments +

government spending + autonomous net export =800+3000+5500+1500=10 800.

b)MPS = 1-MPC = 1-0.95 = 0,05

"M=\frac{1}{MPS}=\frac{1}{0.05}=20"

c)Disposable Income=C+S-Tx=50 232.56-1000-0.3Q=50 232.56-1000-0.3*50 232.56=34 162.79

C=800+0.95Q=800+0.95*50 232.56=48 520.93

S=-800+0.05Q=-800+0.05*50 232.56=1 711.63

d)C=800+0.95Q=800+0.95*50 232.56=48 520.93

e)S=-800+0.05Q=-800+0.05*50 232.56=1 711.63

f)Im = 200 + 0,165Q=200+0.165*50 232.56=8 488.37

g)NX=X-Im=1700-200-0.165Q=1500-0.165Q=1500-8 288.37=-6788.37

h)Government Expenditure=5500+7000=12500

i)Budget Deficit =12 500-1000-0.3Q=12 500-1000-15069.77=-3 569.77

3)

Q=800+0.95Q+3000+5500+1700-(200+0.165Q)

Q=800+0.95q+2500+5500+1700-200-0.165Q

Q-0.785Q=10 300

Q=47 906.98

a)autonomous expenses=800

total autonomous expenses = autonomous consumption + autonomous investments +

government spending + autonomous net export =800+2500+5500+1500=10 300.

b)MPS = 1-MPC = 1-0.95 = 0,05

"M=\frac{1}{MPS}=\frac{1}{0.05}=20"

c)Disposable Income=C+S-Tx=47 906.98-1000-0.3Q=47 906.98-1000-0.3*47 906.98=32 534.89

C=800+0.95Q=800+0.95*47 906.98=46 311.63

S=-800+0.05Q=-800+0.05*47 906.98=1 595.35

d)C=800+0.95Q=800+0.95*47 906.98=46 311.63

e)S=-800+0.05Q=-800+0.05*47 906.98=1 595.35

f)Im = 200 + 0,165Q=200+0.165*47 906.98=8 104.65

g)NX=X-Im=1700-200-0.165Q=1500-0.165Q=1500-7 904.65=-6 404.65

h)Government Expenditure=5500+7000=12500

i)Budget Deficit =12 500-1000-0.3Q=-2 872.09

4)

Q=800+0.95Q+3000+5500+1700-(200+0.165Q)

Q=800+0.95q+2500+5500+1200-200-0.165Q

Q-0.785Q=9 800

0.215Q=9 800

Q=45 581.39

a) autonomous expenses=800

total autonomous expenses = autonomous consumption + autonomous investments +

government spending + autonomous net export =800+2500+5500+1200=10 000.

b)MPS = 1-MPC = 1-0.95 = 0,05

"M=\frac{1}{MPS}=\frac{1}{0.05}=20"

c)Disposable Income=C+S-Tx=45 581.39-1000-0.3Q=45 581.39-1000-0.3*45 581.39=30 906.97

C=800+0.95Q=800+0.95*45 581.39=44 102.32

S=-800+0.05Q=-800+0.05*45 581.39=1 479.07

d)C=800+0.95Q=800+0.95*45 581.39=44 102.32

e)S=-800+0.05Q=-800+0.05*45.581.39=1 479.07

f)Im = 200 + 0,165Q=200+0.165*45.581.39=7 720.93

g)NX=X-Im=1200-200-0.165Q=1000-0.165Q=1000-7 520.93=-6520.93

h)Government Expenditure=5500+7000=12500

i)Budget Deficit =12 500-1000-0.3Q=-2 174.42